OpenStudy (anonymous):

How do I slove this: xy^-1-yx^-1 the division bad sign y^-1-x^-1

OpenStudy (wolf1728):

I'm guessing that you need to solve this? (xy^-1-yx^-1) ÷ (y^-1-x^-1)

OpenStudy (zzr0ck3r):

division bad sign?

OpenStudy (anonymous):

Yeah, I need to solve this! Can you show me how to? Sorry I meant to say division bar sign

OpenStudy (zzr0ck3r):

so \[\frac{xy^{-1}-yx^{-1}}{y^{-1}-x^{-1}}\]?

OpenStudy (anonymous):

Yeah!

OpenStudy (zzr0ck3r):

\[\frac{xy^{-1}-yx^{-1}}{y^{-1}-x^{-1}}=\frac{\frac{x}{y}-\frac{y}{x}}{\frac{1}{y}-\frac{1}{x}}=\frac{\frac{x^2-y^2}{xy}}{\frac{x-y}{xy}}=\frac{x^2-y^2}{xy}*\frac{xy}{x-y}=\frac{x^2-y^2}{x-y}=\\\frac{(x-y)(x+y)}{x-y}=x+y\]as long as\[x\ne y\]

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (zzr0ck3r):

np